OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi divided by five. How do you do this? :/

OpenStudy (anonymous):

\[\pi \div5?\]

OpenStudy (anonymous):

π/5 as in a fraction.

OpenStudy (anonymous):

The easiest way to do this is to begin by drawing a picture.

OpenStudy (anonymous):



OpenStudy (anonymous):

Do you understand that picture? The angle starts at the positive x-axis, goes \(\Large \frac{\pi}{5}\) around the circle and ends there.

OpenStudy (anonymous):

i understand the picture but... what would the angles be? i just see pi/5

OpenStudy (anonymous):

Okay cool, so you understand the picture. Now let's make sure you understand the rest of what you need to know. Do you know what "coterminal" means? Do you know how many radians one whole trip around the circle is?

OpenStudy (anonymous):

coterminal would be the rest of the angle that completes the circle right? and the whole trip would be 2pi?

OpenStudy (anonymous):

A whole trip is 2pi. That's correct. Now, coterminal doesn't mean exactly that. It means that the two angles end at the same spot on the circle. Let's me show you an example of some angles that are all coterminal: